Let f(x) = x + 4 & g(x) = x2 – 5 (fºg)(x)
Answer #1

Solution:

\small f(x)=x+4\textup{ and }g(x)=x^{2}-5

\small (f\circ g)(x)=f(g(x))=g(x)+4=(x^{2}-5)+4=x^{2}-1

\small \boldsymbol{\therefore (f\circ g)(x)=x^{2}-1}

PS: hit the like button